Solve using the multiplication principle first, then the elimination method:
7p+5q=2, 8p-9q=17

Answer by stanbon(75887)

Solve using the multiplication principle first, then the elimination method:
7p+5q=2
8p-9q=17
---------------
Multiply thru 1st by 9.
Multiply thru 2nd by 5.
----
63p + 45q = 18
40p - 45q = 85
------------------
Add to get:
103p = 103
p = 1
----
Solve for "q":
7p + 5q = 2
7 + 5q = 2
5q = -5
q = -1
-----------
Cheers,
Stan H.
========================